The U.S. government is accepting bids to supply MUDDs for the next three years to the military. The contract requires the winner to deliver 4 MUDDs per year. You estimate that the project will require an initial investment in fixed assets of $50,000 which will be depreciated straight line to their expected selling price of $10,000 after the project is finished. Also, a $10,000 expansion in NWC is needed. You estimate that warehouse space which can be leased for $12,000 per year is also needed. Variable cost per MUDD will be $10,000. The firm’s marginal tax rate is 34% and the required return on projects with this level of risk is 15%. What is the minimum price that should be set per MUDD?
Answer #1
1) The minimum price will be one for which NPV = 0.
2) Initial investment:
Fixed assets 50000
Increase in NWC 10000
Total initial investment 60000
3) Annual variable cost (4*10000) 40000
Annual depreciation [(50000-10000)/4] 10000
Lease rent lost 12000
Total before tax expenses 62000
Tax shield at 34% 21080
Annual after tax expenses 40920
Less: Depreciation 10000
Annual after tax cash expenses 30920
4) After tax salvage value = 10000*(1-34%) = 6600
Release of NWC 10000
After tax terminal non-operating cash flow 16600
5) NPV of the above known cash flows = -60000-30920*(1.15^4-1)/(0.15*1.15^4)+16600/1.15^4 = -138785
6) The PV of the after tax sales revenue should be $138785
to get 0 NPV.
7) The after tax annual sales revenue would = 138785*0.15*1.15^4/(1.15^4-1) = $        48,612
Before tax sales revenue = 48612/66% = $        73,654
8) Minimum price to be set per MUDD = 73654/4 = $        18,413
CHECK:
Annual sales revenue = 18413*4 = 73652
Annual variable cost (4*10000) 40000
Annual depreciation [(50000-10000)/4] 10000
Lease rent lost 12000
EBIT 11652
Tax at 34% 3962
NOPAT 7690
Add: Depreciation 10000
Annual OCF 17690
NPV = -60000+17690*(1.15^4-1)/(0.15*1.15^4)+16600/1.15^4 = -4
Almost 0